A plane leaves an airport with a bearing of S41W. After traveling 24 miles, the plane turns and travels on a bearing of N49W for 32 miles. At that time, what's the bearing of the plane from the airport?


1
Expert's answer
2022-06-01T02:54:30-0400

"\angle B=41\degree+49\degree=90\degree," so we have the right triangle ABC.

Let's find AC.

"AC=\sqrt{AB^2+CB^2}=\sqrt{24^2+32^2}=\sqrt{1600}=40."

Now we can calculate "\angle A:"

"sinA=\frac{BC}{AC}=\frac{32}{40}=0.8,"

"\angle A \approx53\degree."




So "\theta=180\degree-(53\degree+41\degree)=86\degree."

 The bearing of the plane from the airport is N86W.


Answer: N86W.